OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (3, 4) and (5, 16)? Help please

OpenStudy (anonymous):

slope,m = y2-y1/x2-x1 = (16-4)/(5-3) = 12/2=6 the equation can be written as, y= mx + c intercept c = y-mx = 4-6*3 = 4-18=-14 equation is y= 6x -14 or 6x-y = 14